OpenStudy (jozelynw):

If the question says insert x=-5 into the equation 4^-x how would you do that? Would you say that 2 negatives equal a positive or just input -5

jimthompson5910 (jim_thompson5910):

you would replace x with -5 so 4^(-x) = 4^(-(-5)) the two negatives cancel to form a positive which is why 4^(-(-5)) = 4^5

OpenStudy (anonymous):

yeah, it equal to positive

OpenStudy (jozelynw):

ok thank you i was just confused

OpenStudy (jozelynw):

what if its a positive number then?

OpenStudy (jozelynw):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

like if x = 9 ?

jimthompson5910 (jim_thompson5910):

if x = 9, then, \[\LARGE 4^{-x} = 4^{-9} = \frac{1}{4^9}\]

OpenStudy (jozelynw):

ok

OpenStudy (jozelynw):

what if x=1

jimthompson5910 (jim_thompson5910):

the negative exponent tells us to take the reciprocal of the base to make the exponent positive \[\LARGE x^{-k} = \frac{1}{x^k}\]

OpenStudy (jozelynw):

so for 2^x+3 what would x=3 be then

jimthompson5910 (jim_thompson5910):

\[\Large 2^x+3\] or \[\Large 2^{x+3}\] ??

OpenStudy (jozelynw):

the second 1

OpenStudy (jozelynw):

i think the answer would be 64

jimthompson5910 (jim_thompson5910):

\[\Large 2^{x+3} = 2^{3+3} = 2^6 = 64\] I agree

OpenStudy (jozelynw):

Ok thanks so much, I'm going to need more help later thou.